# Test-Time Anchoring for Discrete Diffusion Posterior Sampling

### v1

@@ 3.2 Test-Time Anchored Posterior Sampling
In posterior sampling, our goal is to construct a Markov chain whose stationary distribution coincides with the Bayesian posterior: $q({\mathbf{x}}|{\mathbf{y}})\propto q({\mathbf{y}}|{\mathbf{x}})\,q({\mathbf{x}}),$ where ${\mathbf{y}}={\mathcal{A}}({\mathcal{D}}({\mathbf{x}}))+\sigma\varepsilon$ with measurement operator ${\mathcal{A}}(\cdot)$, image decoder ${\mathcal{D}}(\cdot)$, Gaussian noise $\varepsilon\sim\mathcal{N}(0,I)$, and standard deviation $\sigma$. When $\mathcal{A}$} is linear the task reduces to a linear inverse problem; otherwise a nonlinear inverse problem. We approximate $q({\mathbf{x}}|{\mathbf{y}})$ with a tractable sampler $p_{\varphi}({\mathbf{x}}|{\mathbf{y}})$ using only a masked diffusion model $p_{\theta}({\mathbf{x}})$ previously trained to approximate $q({\mathbf{x}})$.
@@
To sample from the posterior $q(\cdot|{\mathbf{y}})$, we construct a Markov chain with the joint distribution defined as: $p_{\varphi}({\mathbf{x}},{\mathbf{z}}_{0:1}|{\mathbf{y}})=p_{\varphi}({\mathbf{z}}_{1}|{\mathbf{y}})\,p_{\varphi}({\mathbf{x}}|{\mathbf{z}}_{0},{\mathbf{y}})\,\prod_{i=1}^{T}p_{\varphi}\!\left({\mathbf{z}}_{s(i)}|{\mathbf{z}}_{t(i)},{\mathbf{y}}\right)$, where ${\mathbf{z}}_{0:1}={\mathbf{z}}_{0},{\mathbf{z}}_{1/T},\dots,{\mathbf{z}}_{1}$. We parameterize measurement conditional transitions by tilting the $\mathbf{y}$}-unconditional transition (4) with the likelihood of measurements given the current estimate:
$p_{\varphi}\!\left({\mathbf{z}}_{s}|{\mathbf{z}}_{t},{\mathbf{y}}\right)\coloneqq\prod_{l=1}^{L}p_{\varphi}({\mathbf{z}}_{s}^{l}|{\mathbf{z}}_{t},{\mathbf{y}}),~\text{where}~p_{\varphi}({\mathbf{z}}_{s}^{l}|{\mathbf{z}}_{t},{\mathbf{y}})\;\coloneqq\;q\!\big({\mathbf{z}}_{s}^{l}|{\mathbf{z}}_{t}^{l},{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t})\big)\,q\!\big({\mathbf{y}}|{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t})\big).$ (6)
Theorem 3.1 (Discrete Diffusion Posterior Sampling (DDPS)).
Given a sample ${\mathbf{x}}\sim q$, let $q(Z_{0:1}|{\mathbf{x}})$ denote the forward noising law of (3). Then, for any measurement ${\mathbf{y}}\sim q(\cdot|{\mathbf{x}})$, the negative log-posterior is bounded by $\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})\leq{\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)\coloneqq\mathbb{E}_{q(Z_{0:1}|{\mathbf{x}})}[-\log p_{\varphi}({\mathbf{x}}|Z_{0})]+$
$\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Big[\frac{\alpha_{t(i)}-\alpha_{s(i)}}{1-\alpha_{t(i)}}\sum_{l=1}^{L}\log\langle{\mathbf{x}}^{l}_{\varphi}(Z_{t(i)}),{\mathbf{x}}^{l}\rangle\bm{1}_{\{Z_{t(i)}^{l}={\mathbf{m}}\}}\Big]-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big].$
Implications. Theorem 3.1 shows that ${\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ is a principled training criterion for discrete posterior samplers. The likelihood-based tilt $\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))$ enforces measurement consistency When $\mathbf{y}$} is absent, the tilting terms vanish and the objective reduces to the standard masked diffusion NELBO (5). The negative cross-entropy term $\log\langle{\mathbf{x}}^{l}_{\varphi}(Z_{t(i)}),{\mathbf{x}}^{l}\rangle$ is zero for revealed tokens and gets supervision for masked tokens, with weights determined by the noise schedule.
For retraining, one can minimize ${\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ with respect to $\varphi$ to obtain a discrete posterior sampler. In practice, however, retraining a large-scale foundation model per task is often infeasible due to excessive compute and lack of training data. We therefore focus on the training-free case.
Theorem 3.2 (Test-time Anchored Posterior Sampling (APS)).
Given a sample ${\mathbf{x}}\sim q$, let $q(Z_{0:1}|{\mathbf{x}})$ denote the forward noising law of (3). Suppose the pretrained network $p_{\theta}({\mathbf{x}})$ closely approximates the unconditional prior $q({\mathbf{x}})$. Then, for any measurement ${\mathbf{y}}\sim q(\cdot|{\mathbf{x}})$, the negative log-posterior is bounded by $\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})\leq{\mathcal{L}}_{\mathrm{APS}}({\mathbf{x}},{\mathbf{y}};\varphi)\coloneqq{\mathcal{L}}_{\mathrm{NELBO}}({\mathbf{x}};\theta)+$
$\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Big[\frac{\alpha_{s(i)}-\alpha_{t(i)}}{1-\alpha_{t(i)}}\,\sum_{l=1}^{L}\log\frac{\langle{\mathbf{x}}^{l}_{\theta}(Z_{t(i)}),{\mathbf{x}}^{l}\rangle}{\langle{\mathbf{x}}^{l}_{\varphi}(Z_{t(i)}),{\mathbf{x}}^{l}\rangle}\,\bm{1}_{\{Z^{l}_{t(i)}={\mathbf{m}}\}}\Big]-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big].$
Implications. Theorem 3.2 shows that posterior sampling can be performed without retraining by reusing a pretrained masked diffusion model.
  • Efficient test-time training. The bound ${\mathcal{L}}_{\mathrm{APS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ is expressed in terms of the pretrained NELBO from (5). Since this term is constant with respect to the new parameters $\varphi$, it can be ignored during optimization. As a result, test-time training only needs to update the lightweight adaptation and measurement-consistency terms, while reusing the fixed pretrained network ${\mathbf{x}}_{\theta}(\cdot)$. This avoids backpropagation through the large-scale denoiser (e.g., billions of parameters), making posterior sampling feasible and efficient at test time.
  • Training-free inference. Although efficient test-time training requires paired $({\mathbf{x}},{\mathbf{y}})$ data, in posterior sampling we only observe $\mathbf{y}$}. Interestingly, the bound ${\mathcal{L}}_{\mathrm{APS}}$ points to training-free inference by substituting the pretrained model prediction ${\mathbf{x}}^{l}_{\theta}(Z_{t(i)})$ in place of ${\mathbf{x}}^{l}$ (see §3.2.1). This makes posterior sampling feasible directly from measurements, without retraining or labeled pairs.
  • Adaptation gap. The log-ratio terms capture the mismatch between unconditional predictions ${\mathbf{x}}_{\theta}(Z_{t(i)})$ and adapted posterior predictions ${\mathbf{x}}_{\varphi}(Z_{t(i)})$, active only at masked positions.
  • Measurement consistency. The final summation enforces agreement with the measurement likelihood, ensuring the sampler produces samples consistent with observed measurements $\mathbf{y}$}.
Discussion. In summary, retraining a new network $\varphi$ for every task would require backpropagation through massive denoisers (e.g., 8B parameters in MMaDA), which is computationally prohibitive. Our theoretical results show that posterior sampling can be done efficiently by reusing the pretrained model and optimizing only lightweight parameters at test time. Next, we introduce two key ingredients—Quantized Expectation3.2.1) and Anchored Remasking3.2.2)—that make this training-free posterior sampling practically implementable. These two ideas together form our Algorithm 1: Anchored Posterior Sampling (APS); please see Appendix B.2 for a detailed discussion.
3.2.1 Quantized Expectation
We propose quantized expectation to compute $\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))$ in a differentiable manner. Building on the training-free inference implication of Theorem 3.2, we note that in posterior sampling only the measurements $\mathbf{y}$} are observed. Hence, we replace the ground-truth token ${\mathbf{x}}^{l}$ with the token mixture predicted by the pretrained model ${\mathbf{x}}^{l}_{\theta}(Z_{t(i)})$, which simplifies the upper bound to
$\hat{{\mathcal{L}}}_{\mathrm{APS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ $=\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Bigg[\frac{\alpha_{t(i)}-\alpha_{s(i)}}{1-\alpha_{t(i)}}\sum_{l=1}^{L}\log\langle{\mathbf{x}}^{l}_{\varphi}(Z_{t(i)}),{\mathbf{x}}^{l}_{\theta}(Z_{t(i)})\rangle\,\bm{1}_{\{Z^{l}_{t(i)}={\mathbf{m}}\}}\Bigg]$
$\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big]+\text{constants}.$ (7)
The measurement likelihood takes the form $q({\mathbf{y}}|{\mathbf{x}})\propto\exp(-\tfrac{\|{\mathbf{y}}-{\mathcal{A}}({\mathcal{D}}({\mathbf{x}}))\|_{2}^{2}}{2\sigma^{2}}),$ where ${\mathcal{D}}(\cdot)$ denotes the image decoder. The denoiser outputs token probabilities ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)})=\{{\mathbf{x}}^{l}_{\varphi}({\mathbf{z}}_{t(i)})\}_{l=1}^{L}$. Since $\mathcal{D}$} takes a sequence of discrete tokens (or their one-hot encodings) as input, a naive approach would sample ${\mathbf{x}}=({\mathbf{x}}^{1},\dots,{\mathbf{x}}^{L})\sim{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)})$ and compute $q({\mathbf{y}}|{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)}))\propto\exp(-\tfrac{\|{\mathbf{y}}-{\mathcal{A}}({\mathcal{D}}({\mathbf{x}}))\|_{2}^{2}}{2\sigma^{2}}).$ This introduces sampling noise and also makes $\hat{{\mathcal{L}}}_{\mathrm{APS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ non-differentiable. Alternatively, one could update $\varphi$ by employing policy gradient rule for non-differentiable rewards as in SVDD-PM (Li et al., 2024). However, this leads to sparse updates and inferior sample quality as shown in Table 1.
@@ 3.2.1 Quantized Expectation
Differentiability has propelled posterior sampling to achieve state-of-the-art results using continuous diffusion (Chung et al., 2023; Rout et al., 2023; 2024; Zhang et al., 2025). To restore differentiability in discrete diffusion, we parameterize ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)})=\mathrm{Softmax}(\varphi_{t(i)})$ with $\varphi_{t(i)}=\{\varphi^{l}_{t(i)}\}_{l=1}^{L}\in\mathbb{R}^{K\times L}$ containing logits $\varphi^{l}_{t(i)}$ over the codebook $\{{\mathbf{c}}_{k}\in{\mathcal{C}}\}$ for each position $l$. We compute the expected embedding $\bar{{\mathbf{x}}}^{l}=\sum_{k=1}^{K}{\mathbf{c}}_{k}\,{\mathbf{x}}^{l}_{\varphi}({\mathbf{z}}_{t(i)})\in\mathbb{R}^{d},$ and then quantize it using LFQ (2(Yu et al., 2024): ${\mathbf{x}}^{l}={\mathcal{Q}}_{\mathrm{lfq}}(\bar{{\mathbf{x}}}^{l})\in\{-1,+1\}^{d}.$ We then apply the straight-through estimator (Van Den Oord et al., 2017) to obtain an image $\hat{{\mathbf{x}}}={\mathcal{D}}(\tilde{{\mathbf{x}}}),~\tilde{{\mathbf{x}}}=\bar{{\mathbf{x}}}+\big[{\mathbf{x}}-\bar{{\mathbf{x}}}\big]_{\mathrm{sg}},$ where $\mathrm{sg}$ denotes the stop-gradient operator. Finally, we compute the differentiable likelihood as $q({\mathbf{y}}|{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)}))\propto\exp(-\tfrac{\|{\mathbf{y}}-{\mathcal{A}}(\hat{{\mathbf{x}}})\|_{2}^{2}}{2\sigma^{2}}).$
Discussion. We minimize the objective (3.2.1) at every step $t$ to obtain the optimal logits $\varphi^{*}_{t(i)}$. This ensures that the optimized probabilities ${\mathbf{x}}_{\varphi^{*}}({\mathbf{z}}_{t(i)})$ remain close to the prior predictions ${\mathbf{x}}_{\theta}({\mathbf{z}}_{t(i)})$ while being adapted to the measurements. To the best of our knowledge, this paper takes the first step towards quantizing the expected codebook entry for posterior sampling in discrete diffusion. This allows gradients from the measurement loss to propagate through $\bar{{\mathbf{x}}}$ and update the entire conditional probability table ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)})$, strengthening gradient-based guidance and accurate posterior sampling.
3.2.2 Anchored Remasking
In masked diffusion, the reverse process progressively unmasks tokens. Most discrete samplers (Austin et al., 2021; Chang et al., 2022; Lou et al., 2024; Sahoo et al., 2024; Shi et al., 2024) choose to unmask tokens in random order. ADLM (Rout et al., 2025a) shows that prioritizing “anchor” tokens (e.g., nouns or verbs in language, rather than articles or conjunctions) early in the reverse process reduces conditional entropy of the remaining sequence and improves generation quality. To decode anchor tokens, ADLM jointly trains an anchor network in addition to the standard denoising network using an anchored NELBO objective. Distinct from ADLM, we propose a training-free variant of anchored denoising, enabling posterior sampling with a pretrained model.
Let $\varphi$ be a minimizer of (3.2.1), resulting in ${\mathbf{x}}^{*}_{\varphi}({\mathbf{z}}_{t})=\{({\mathbf{x}}^{*}_{\varphi}({\mathbf{z}}_{t}))^{l}\}_{l=1}^{L}$ that represents categorical distributions over tokens at each position $l\in\{1,\dots,L\}$ at time step $t$. Anchored remasking selects a subset of positions ${\mathcal{P}}_{t}\subseteq\{1,\dots,L\}$ to decode early. The selection is based on the confidence of quantized tokens $\mathbf{x}$} (as defined in §3.2.1) under the posterior estimate ${\mathbf{x}}^{*}_{\varphi}({\mathbf{z}}_{t})$. Importantly, the posterior estimate is a function of the $L$-length sequence ${\mathbf{z}}_{t}$, and hence encodes the joint relation across all tokens; thus, anchored remasking is a function of all tokens, as compared to standard remasking (Chang et al., 2022; Yang et al., 2025) that depend only on per-token confidence. Formally, we compute the confidence of ${\mathbf{x}}^{l}$ as $\kappa^{l}_{t}=\langle({\mathbf{x}}^{*}_{\varphi}({\mathbf{z}}_{t}))^{l},{\mathbf{x}}^{l}\rangle,$ and choose anchor positions as ${\mathcal{P}}_{t}=\{\,l:\kappa^{l}_{t}\geq\tau_{t}\,\},$ where $\tau_{t}$ is an adaptive threshold based on the cosine schedule from MMaDA. We then update the state by fixing anchor tokens and remasking the rest: ${\mathbf{z}}_{s}^{l}={\mathbf{x}}$ with probability $\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}$ if $l\in{\mathcal{P}}_{t}$ else $\mathbf{m}$}. Once a token is unmasked, it remains fixed in subsequent steps.
Discussion. Diffusion language models tend to be overconfident on low-information tokens such as articles (“a”, “an” or “the”) or conjunctions (Rout et al., 2025a); similarly, discrete image samplers using independent per-token confidence often unmask background pixels first. In contrast, our method leverages the joint posterior ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t})$ to identify anchor tokens consistent with the measurements. This leads to earlier decoding of informative tokens (e.g., a bird against a flat background), enabling faster reconstruction and improved likelihood of generated samples; refer §B.3 for details.

### v2

@@ 3 Test-Time Anchored Posterior Sampling
In posterior sampling, our goal is to construct a Markov chain to sample from the posterior: $q({\mathbf{x}}|{\mathbf{y}})\propto q({\mathbf{y}}|{\mathbf{x}})\,q({\mathbf{x}}),$ where ${\mathbf{y}}={\mathcal{A}}({\mathcal{D}}({\mathbf{x}}))+\sigma\varepsilon$ with measurement operator ${\mathcal{A}}(\cdot)$, image decoder ${\mathcal{D}}(\cdot)$, Gaussian noise $\varepsilon\sim\mathcal{N}(0,I)$, and standard deviation $\sigma$. When $\mathcal{A}$} is linear the task reduces to a linear inverse problem; otherwise a nonlinear inverse problem. We approximate $q({\mathbf{x}}|{\mathbf{y}})$ with a tractable sampler $p_{\varphi}({\mathbf{x}}|{\mathbf{y}})$ using a masked diffusion model $p_{\theta}({\mathbf{x}})$ previously trained to approximate $q({\mathbf{x}})$.
@@
To sample from the posterior $q(\cdot|{\mathbf{y}})$, we construct a Markov chain with the joint distribution defined as: $p_{\varphi}({\mathbf{x}},{\mathbf{z}}_{0:1}|{\mathbf{y}})=p_{\varphi}({\mathbf{z}}_{1}|{\mathbf{y}})\,p_{\varphi}({\mathbf{x}}|{\mathbf{z}}_{0},{\mathbf{y}})\,\prod_{i=1}^{T}p_{\varphi}\!\left({\mathbf{z}}_{s(i)}|{\mathbf{z}}_{t(i)},{\mathbf{y}}\right)$, where ${\mathbf{z}}_{0:1}={\mathbf{z}}_{0},{\mathbf{z}}_{1/T},\dots,{\mathbf{z}}_{1}$. We parameterize measurement conditional transitions by tilting the unconditional transition (4) with the likelihood of measurements given the current estimate: $p_{\varphi}\!\left({\mathbf{z}}_{s}|{\mathbf{z}}_{t},{\mathbf{y}}\right)\coloneqq\prod_{l=1}^{L}p_{\varphi}({\mathbf{z}}_{s}^{l}|{\mathbf{z}}_{t},{\mathbf{y}}),\text{where}$
$p_{\varphi}({\mathbf{z}}_{s}^{l}|{\mathbf{z}}_{t},{\mathbf{y}})\;{\propto}\;q\!\big({\mathbf{z}}_{s}^{l}|{\mathbf{z}}_{t}^{l},{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t})\big)\,q\!\big({\mathbf{y}}|{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t}{;{\mathbf{z}}_{s}^{l}})\big).$ (6)
We can compute the measurement likelihood $q({\mathbf{y}}|{\mathbf{x}})$ given a sequence of tokens ${\mathbf{x}}=({\mathbf{x}}^{1},{\mathbf{x}}^{2},\dots,{\mathbf{x}}^{L})$, where each ${\mathbf{x}}^{l}$ is a one-hot vector. During inference, we have access to the model output ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t})$ rather than the ground-truth sequence $\mathbf{x}$}. We define the log likelihood given the model’s output as $\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t}))\coloneqq\mathbb{E}_{{\mathbf{x}}\sim{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t})}[\,\log q({\mathbf{y}}|{\mathbf{x}})\,]$. Similarly, $\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t};{\mathbf{z}}_{s}^{l}))$ denotes the same log likelihood, but with ${\mathbf{x}}^{l}$ replaced by ${\mathbf{z}}_{s}^{l}$ while keeping all other positions drawn from ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t})$. We denote the normalizing factor in (6) by $\mathcal{Z}^{l}_{\varphi}({\mathbf{z}}_{t},{\mathbf{y}})$.
@@
Theorem 3.1 (Discrete Diffusion Posterior Sampling (DDPS)).
Given a sample ${\mathbf{x}}\sim q$, let $q(Z_{0:1}|{\mathbf{x}})$ denote the forward noising law of (3). Then, for any measurement ${\mathbf{y}}\sim q(\cdot|{\mathbf{x}})$, $\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})\leq{\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)\coloneqq\mathbb{E}_{q(Z_{0}|{\mathbf{x}})}[-\log p_{\varphi}({\mathbf{x}}|Z_{0})]+{\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t}|{\mathbf{x}})}\big[\sum_{l=1}^{L}\log\mathcal{Z}^{l}_{\varphi}(Z_{t},\mathbf{y})\big]+}$ $\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t}|{\mathbf{x}})}\Bigg[\frac{\alpha_{t}-\alpha_{s}}{1-\alpha_{t}}\sum_{l=1}^{L}\log\langle{\mathbf{x}}^{l}_{\varphi}(Z_{t}),{\mathbf{x}}^{l}\rangle\bm{1}_{\{Z_{t}^{l}={\mathbf{m}}\}}\\$ $\mathbb{E}_{q({Z_{s}}|Z_{t},{\mathbf{x}})}\Big[\sum_{l=1}^{L}\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t};Z_{s}^{l}))\Big]\Bigg].$
Implication 3.2.
Theorem 3.1 shows that ${\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ is a principled training criterion for discrete posterior samplers. The likelihood-based tilt $\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)};Z_{s}))$ enforces measurement consistency. When $\mathbf{y}$} is absent, the tilting term vanishes and the objective reduces to the standard masked diffusion NELBO (5). We can treat $\mathcal{Z}_{\varphi}$ as a stop-gradient term—used to normalize the probability distribution (e.g., via Softmax) but excluded during gradient computation (Sohl-Dickstein et al., 2015; Dhariwal & Nichol, 2021; Ho & Salimans, 2022; Murata et al., 2024). Although expensive, a more formal treatment for the gradient of the log partition function can be obtained through importance sampling (Kahn, 1950), Hamiltonian Monte Carlo with leapfrog steps (MacKay, 2003), or noise contrastive estimation (Gutmann & Hyvärinen, 2010). The cross-entropy term $\log\langle{\mathbf{x}}^{l}_{\varphi}(Z_{t}),{\mathbf{x}}^{l}\rangle$ gets supervision through the masked tokens, with weights determined by the noise schedule.
For retraining, one can minimize ${\mathcal{L}}_{\mathrm{DDPS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ with respect to $\varphi$ to obtain a discrete posterior sampler. In practice, however, retraining a large-scale foundation model per task is often expensive due to excessive compute and lack of training data. We therefore focus on the training-free case.
Theorem 3.3 (Test-time Anchored Posterior Sampling (APS)).
Given a sample ${\mathbf{x}}\sim q$, let $q(Z_{0:1}|{\mathbf{x}})$ denote the forward noising law of (3). Suppose the pretrained network $p_{\theta}({\mathbf{x}})$ closely approximates the unconditional prior $q({\mathbf{x}})$. Then, for any measurement ${\mathbf{y}}\sim q(\cdot|{\mathbf{x}})$, the negative log-posterior $\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})\leq{\mathcal{L}}_{\mathrm{APS}}({\mathbf{x}},{\mathbf{y}};\varphi)\coloneqq{\mathcal{L}}_{\mathrm{NELBO}}({\mathbf{x}};\theta){\;+\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t}|{\mathbf{x}})}\Big[\sum_{l=1}^{L}\log\mathcal{Z}^{l}_{\varphi}(Z_{t},\mathbf{y})\Big]}+$ $\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t}|{\mathbf{x}})}\Bigg[\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}\,\sum_{l=1}^{L}\log\frac{\langle{\mathbf{x}}^{l}_{\theta}(Z_{t}),{\mathbf{x}}^{l}\rangle}{\langle{\mathbf{x}}^{l}_{\varphi}(Z_{t}),{\mathbf{x}}^{l}\rangle}\,\bm{1}_{\{Z^{l}_{t}={\mathbf{m}}\}}\\$ $\hskip 17.22217pt-\mathbb{E}_{q({Z_{s}}|Z_{t},{\mathbf{x}})}\Big[\sum_{l=1}^{L}\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t}{;Z_{s}^{l}}))\Big]\Bigg].$
Implication 3.4.
Theorem 3.3 shows posterior sampling without additional training by reusing a pretrained model.
  • Efficient test-time training. The ${\mathcal{L}}_{\mathrm{APS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ bound is expressed in terms of NELBO (5). Since this term is constant with respect to the new parameters $\varphi$, it can be ignored during optimization. The normalizing constant is treated as in Implication 3.2. As a result, test-time training only needs to update the lightweight adaptation and measurement-consistency terms, while reusing the fixed pretrained network ${\mathbf{x}}_{\theta}(\cdot)$. This avoids backpropagation through the large denoiser (e.g., billions of parameters), making posterior sampling efficient at test time.
  • Training-free inference. Although test-time training requires paired $({\mathbf{x}},{\mathbf{y}})$ data, in posterior sampling we only observe $\mathbf{y}$}. Importantly, the bound ${\mathcal{L}}_{\mathrm{APS}}$ motivates a surrogate optimization objective rather than an exact sampling procedure (see §3.1), placing APS in the class of variationally guided diffusion samplers rather than asymptotically unbiased MCMC methods.
  • Adaptation gap. The log-ratio terms capture the mismatch between unconditional predictions ${\mathbf{x}}_{\theta}(Z_{t(i)})$ and adapted posterior ${\mathbf{x}}_{\varphi}(Z_{t(i)})$, active only at masked positions.
  • Measurement consistency. The final summation ensures sample consistency with the observed measurements $\mathbf{y}$}.
Next, we introduce two approximations: Quantized Expectation3.1) and Anchored Remasking3.2) that make our approach practically implementable. These two ideas together form our Algorithm 1: Anchored Posterior Sampling (APS); please see Appendix C.2 for a detailed discussion.